3log8x + 2log8x = 8
How do I solve?

Answer by Boreal(15235)

first make it log (8x)^3+log (8x)^2=8
Then can make it log{(8x)^3*(8x)^2]=8, since the addition separately is multiplication under one log.
Then raise everything to the 10 power, and 10^8=100000000
8^5*x^5=10000000
divide by 8^5 to get 3051.76, don't round, then raise both sides to the (1/5) power
x=4.976
3 log (39.81, when not rounded)=4.8
2 log (39.81, when not rounded)=3.2